OpenStudy (anonymous):

find the area of the shaped reigion r=root 0

OpenStudy (anonymous):

find the area of the shaped region r=root θ

OpenStudy (anonymous):

find area of circle divide it by 4 to get area of one portion total area - area of one portion

OpenStudy (isaiah.feynman):

@sangya21 It doesn't work that way. That's a spiral.

OpenStudy (anonymous):

Sorry I thought thats a circle.

OpenStudy (anonymous):

The area's given by \[\int_{\pi/2}^{2\pi}\int_0^{\sqrt\theta} r~dr~d\theta=\frac{1}{2}\int_{\pi/2}^{2\pi}\theta~d\theta\]